OpenStudy (anonymous):

Graping y=x ^ 2 / (2 – x) usin derivatives .Plz Help!

OpenStudy (anonymous):

Thanks Ishan. :)

OpenStudy (anonymous):

I found the vertcal asymptote to be x=2, no horz asympte, and y=-x oblique asymptote

OpenStudy (anonymous):

I lso found (0,0) to be a pont on the graph

OpenStudy (anonymous):

Hmm I haven't done this before, I will see if I can get it. \[y' = \frac{2x(2-x)+x^2}{(2-x)^2}\] \[y'=0\implies \frac{4x-x^2}{(2-x)^2} = 0\]Yes, vertical asymptote at x=2. And minima at 0.

OpenStudy (anonymous):

The problem is how can we have a asmpytote of y=-x,when 0,0 id a point on the graph

OpenStudy (anonymous):

Never mind the last sentence

OpenStudy (anonymous):

actually y=-x-2 is an oblique asymptote

OpenStudy (anonymous):

asymptotes can be crossed... only the vertical asymptotes can never be crossed.

OpenStudy (anonymous):

I have found everything,including the second derivative ,concavity and increading/decreasing points

OpenStudy (anonymous):

but cant graph it

OpenStudy (anonymous):

Can you help me graph it

OpenStudy (anonymous):

From (0,2). Second derivative \(\large\frac{-8}{(2-x)^3}\).

OpenStudy (anonymous):

:o Where is my drawing??? :(

OpenStudy (anonymous):

:( I got a graph from (-2,2] as f''(x) < 0.